OpenStudy (anonymous):

Together, Louisa and Jill scored more than 30 points in the basketball game. If Jill scored 12 points, how many points P did Louisa score? A. P > 18 B. P > 42 C. P < 42 D. P < 18

OpenStudy (anonymous):

i think it is (D)

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

if P<18 and jill scored 12 points... let's say P scored 17 (that's less than 18 after all) so their total will be 12 + 17 = 29 <--this is NOT greater than 30

OpenStudy (anonymous):

(C)

OpenStudy (anonymous):

so less than 42

OpenStudy (lgbasallote):

if P is 43 then 12 + 43 = 55 <--still qualifies as more than 30

OpenStudy (lgbasallote):

@Hero perhaps it's wiser not to use equal signs in inequalities?

hero (hero):

I meant to use an inequality sign

hero (hero):

l + j > 30 j = 12 l + 12 > 30 l > 30 - 12 l > ?